OpenStudy (anonymous):

Use the addition formulas for sine and cosine to simplify the expression. cos(6u)cos(7u)+sin(6u)sin(7u)

myininaya (myininaya):

cos(7u-6u)=cos(1u)=cos(u)

OpenStudy (anonymous):

what happen to sin?

myininaya (myininaya):

you know that cos(x-y)=cosxcosy-sinxsiny right?